# Sparse Covariance Neural Networks

### v1

@@ 6 Related Works
Sparse PCA. The instability of the finite-sample covariance matrix is widely studied for PCA (Jolliffe, 2002; Paul, 2007; Baik et al., 2005) and various regularized estimators have been proposed including covariance shrinkage (Ledoit & Wolf, 2003; 2012), lasso penalties (Bien & Tibshirani, 2011; Han Liu & Zhao, 2014), and thresholding (Bickel & Levina, 2008b; Deshpande & Montanari, 2016). We leverage here hard and soft thresholding given their benefits in PCA in low-data sparse-covariance settings and study their impact on VNN stability.
@@
Stability of GNNs-VNNs. VNNs can be seen as graph convolutional neural networks operating on the covariance graph (Sihag et al., 2022), which draws analogies with PCA, but have been shown to be more stable extending the small perturbation analysis of GNNs (Gama et al., 2020b; Kenlay et al., 2021b; a; Levie et al., 2021; Maskey et al., 2023; Gao et al., 2023) to covariances. Due to their robustness, they have proved successful in brain data processing (Sihag et al., 2024), transferability to large-dimension data (Sihag et al., 2023), temporal settings (Cavallo et al., 2024b) and biased datasets (Cavallo et al., 2024a). All these results, however, hold for the dense covariance matrix which is suboptimal in low-data regimes and computationally heavy. Here, we propose S-VNNs and study their stability when the true covariance matrix is sparse and dense. This merits a different treatment than the perturbation assumption studied in (Sihag et al., 2022) and overcomes the limitations of dense VNNs. Our stochastically sparsified VNN draws analogies with the stability of GNNs to random link drops (Gao et al., 2021a; b) but we generalize those findings to the more challenging case where all sparsified probabilities differ rather than being identical. We also propose design strategies for the sparsified probabilities based on our theoretical analysis.
@@
Graph sparsification in GNNs. Graph sparsification techniques have been used for multiple purposes in GNNs but predominantly to improve: (i) scalability in large and dense graphs (Hamilton et al., 2018; Zhang et al., 2019; Zeng et al., 2021; Peng et al., 2022; Srinivasa et al., 2020; Ye & Ji, 2021); (ii) GNN expressiveness in graph classification tasks and alleviate oversmoothing (Papp et al., 2021; Rong et al., 2020; Fang et al., 2023; Zheng et al., 2020; Morris et al., 2021); and (iii) enhance interpretability (Rathee et al., 2021; Li et al., 2023; Naber et al., 2024). Our S-VNNs aim on the one hand to improve scalability and on the other to preserve the tractability and theoretical links with covariance-based data processing. While our sparsification approaches can be applied also to weighted graphs (see Appendix H), in this work we focus on covariance sparsification to characterize theoretically how this affects stability w.r.t. covariance estimation errors and spurious correlations.
7 Conclusion
@@
In this work, we develop sparse covariance neural networks (S-VNNs) and study the effects of sparsification techniques on covariance neural networks. We show that S-VNNs are more stable to the finite-data sample effect and more computationally efficient. In particular, when the true covariance is sparse, we propose S-VNNs with hard/soft thresholding and show that they transfer to the nominal VNN with a rate inversely proportional to the square root of the number of data points. When the true covariance is dense, we put forth a stochastic sparsified approach to reduce the computational complexity while maintaining mathematical tractability and stability. In the framework of stochastic sparsification, we propose principled strategies, i.e., ACV and RCV, that align with the theoretical observations to tune sparsification impact and computation efficiency. Experimental results on four real datasets show that the proposed S-VNNs improve substantially on the computation time and achieve competitive or better performance compared with nominal VNNs, thus validating that spurious correlations are present and damaging in real datasets. Although VNNs can be applied to several other regularized covariance estimators (Ledoit & Wolf, 2003; Bickel & Levina, 2008a; Bien & Tibshirani, 2011; Friedman et al., 2007) to reduce the impact of spurious correlations, we here focus on thresholding since, unlike other estimators, it allows to control the sparsity of the covariance matrix (fundamental for time and memory efficiency), and allows for theoretical tractability. Future work will investigate other covariance estimators and characterize the regularization effect of stochastic sparsification, which allows the model to observe multiple instances of the same covariance with potential benefits for robustness, expressiveness and reduced oversmoothing.
Reproducibility statement. The proofs of all theoretical results are collected in the appendix. For experimental results, we make the code available at the repository https://github.com/andrea-cavallo-98/Sparse-VNN and we provide references to all the datasets in the following. The preprocessing functions for all datasets are made available in the code.
We report additional experimental details for reproducibility in Sec. I.2.
Acknowledgements. Part of this work was supported by the TU Delft AI Labs program, the NWO OTP GraSPA proposal #19497, and the NWO VENI proposal 222.032.
Appendix A Background on PCA and Connection with VNNs
PCA is a change of basis that aims at minimizing the redundancy of the transformed data while maintaining their statistical properties of interest (Shlens, 2014; Jolliffe, 2002). PCA achieves this by projecting the data onto a space where the covariance matrix of the transformed data is diagonal, i.e., the off-diagonal elements, which characterize the redundancy of the data, are all zeros. More formally, given a dataset $\mathbf{X}\in\mathbb{R}^{N\times t}$, where each column is a data sample $\mathbf{x}\in\mathbb{R}^{N}$, PCA applies a transformation $\mathbf{\tilde{X}}=\mathbf{T}^{\mathsf{T}}\mathbf{X}$ with $\mathbf{T}\in\mathbb{R}^{N\times N}$ such that the covariance of the transformed data $\mathbf{\tilde{X}}\mathbf{\tilde{X}}^{\mathsf{T}}$ is diagonal. This goal is achieved by setting $\mathbf{T}=\mathbf{\hat{V}}$, where $\mathbf{\hat{C}}=\mathbf{X}\mathbf{X}^{\mathsf{T}}=\mathbf{\hat{V}}\mathbf{\hat{\Lambda}}\mathbf{\hat{V}}^{\mathsf{T}}$ is the eigendecomposition of the covariance matrix of the data, since $\mathbf{\tilde{X}}\mathbf{\tilde{X}}^{\mathsf{T}}=\mathbf{\hat{V}}^{\mathsf{T}}\mathbf{X}\mathbf{X}^{\mathsf{T}}\mathbf{\hat{V}}=\mathbf{\hat{V}}^{\mathsf{T}}(\mathbf{\hat{V}}\mathbf{\hat{\Lambda}}\mathbf{\hat{V}}^{\mathsf{T}})\mathbf{\hat{V}}=\mathbf{\hat{\Lambda}}$ due to the orthonormality of the covariance eigenvectors. PCA is also commonly used for dimensionality reduction by selecting only some of the eigenvectors of the covariance matrix to define the new basis, and several techniques exist to select the number of components, such as looking at the eigenvalues of the covariance and selecting the number corresponding to the largest eigenvalue drop. PCA is equivalent to the graph Fourier transform (GFT) for a graph defined by the covariance $\mathbf{\hat{C}}$ (Gama et al., 2020c; Isufi et al., 2024). Indeed, the GFT projects data on the space defined by the eigenvectors of the graph shift operator $\mathbf{S}\in\mathbb{R}^{N\times N}$ (a matrix describing the graph structure, e.g., adjacency or Laplacian), i.e., $\mathbf{\tilde{X}}=\mathbf{U}^{\mathsf{T}}\mathbf{X}$ given the eigendecomposition $\mathbf{S}=\mathbf{U}\mathbf{W}\mathbf{U}^{\mathsf{T}}$. If we consider the graph described by the sample covariance matrix, i.e., $\mathbf{S}=\mathbf{\hat{C}}$, then PCA and GFT perform the same operation.
The connection between PCA and VNNs emerges by looking at the action of the covariance filter in the spectral domain. For a data sample $\mathbf{x}$, taking the GFT of the covariance filter leads to $\mathbf{\tilde{u}}=\mathbf{\hat{V}}^{\mathsf{T}}\mathbf{u}=\mathbf{\hat{V}}^{\mathsf{T}}\sum_{k=0}^{K}h_{k}[\mathbf{\hat{V}}\mathbf{\hat{\Lambda}}\mathbf{\hat{V}}^{\mathsf{T}}]^{k}\mathbf{x}=\sum_{k=0}^{K}h_{k}\mathbf{\hat{\Lambda}}^{k}\mathbf{\hat{V}}^{\mathsf{T}}\mathbf{x}$, equivalent to $\sum_{k=0}^{K}h_{k}\hat{\lambda}_{i}^{k}\mathbf{\tilde{x}}_{i}=h(\hat{\lambda}_{i})\mathbf{\tilde{x}}_{i}$ for the $i$-th entry; that is, the filter is a polynomial in the covariance eigenvalues that modulates each component of the GFT (i.e., PCA) of the signal $\mathbf{x}$ separately (cf. Sec. 3). This shows that the covariance filter processes the principal components of the data, and there exists a set of coefficients $h_{k}$ such that a covariance filter performs the same operation as PCA (Sihag et al., 2022, Theorem 1).
Appendix B Proof of Lemma 1
@@
$\begin{split}\left\|\mathbf{V}^{\mathsf{T}}\mathbf{x}-\mathbf{\hat{V}}^{\mathsf{T}}\mathbf{x}\right\|&=\left\|\sum_{i=1}^{N}(\mathbf{v}_{i}-\mathbf{\hat{v}}_{i})^{\mathsf{T}}x_{i}\right\|\leq\sum_{i=1}^{N}\|\mathbf{v}_{i}-\mathbf{\hat{v}}_{i}\||x_{i}|\\ &\leq\sum_{i=1}^{N}\|\mathbf{v}_{i}-\mathbf{\hat{v}}_{i}\|\leq N\max_{i}\left\|\mathbf{v}_{i}-\mathbf{\hat{v}}_{i}\right\|.\end{split}$ (3)
where we used triangle inequality, the fact that $\|\mathbf{x}\|\leq 1$ and the fact that $\|(\mathbf{v}_{i}-\mathbf{\hat{v}}_{i})^{\mathsf{T}}\|=\|\mathbf{v}_{i}-\mathbf{\hat{v}}_{i}\|$. Using the sin-theta theorem (Stewart & guang Sun, 1990, Theorem V.3.6) (cf. equation 18), we have that
$\max_{i}\left\|\mathbf{v}_{i}-\mathbf{\hat{v}}_{i}\right\|\leq\frac{\sqrt{2}\|\mathbf{E}\|}{\min_{i}|\lambda_{i}-\lambda_{i+1}|}$ (4)
where $\|\mathbf{E}\|=\|\mathbf{C}-\mathbf{\hat{C}}\|$. We now leverage the result from (Vershynin, 2018, Theorem 5.6.1) which shows that $\|\mathbf{E}\|\leq\mathcal{O}(t^{-1/2})$ with high probability, and this concludes the proof. ∎
Appendix C Proof of Thms. 2 and 3
C.1 Preliminaries
@@
Consider the true and the thresholded sample covariance matrices and their respective eigendecompositions $\mathbf{C}=\mathbf{V}\mathbf{\Lambda}\mathbf{V}^{\mathsf{T}}$ and $\mathbf{\bar{C}}=\mathbf{\bar{V}}\bar{\mathbf{\Lambda}}\mathbf{\bar{V}}^{\mathsf{T}}$. The following holds:
$\frac{1}{\sqrt{2}}|\mathbf{v}_{j}^{\mathsf{T}}\mathbf{\bar{v}}_{i}|\leq\sqrt{1-|\mathbf{v}_{j}^{\mathsf{T}}\mathbf{\bar{v}}_{j}|}\quad\forall i\neq j.$ (5)
@@
We first observe that, due to the orthonormality of $\mathbf{V}$ and $\mathbf{\bar{V}}$, we have
$\frac{1}{2}\|\mathbf{v}_{j}-\mathbf{\bar{v}}_{j}\|^{2}=\frac{1}{2}(\mathbf{v}_{j}-\mathbf{\bar{v}}_{j})^{\mathsf{T}}(\mathbf{v}_{j}-\mathbf{\bar{v}}_{j})=$ (6)
$\frac{1}{2}(\mathbf{v}_{j}^{\mathsf{T}}\mathbf{v}_{j}-2\mathbf{v}_{j}^{\mathsf{T}}\mathbf{\bar{v}}_{j}+\mathbf{\bar{v}}_{j}^{\mathsf{T}}\mathbf{\bar{v}}_{j})=1-|\mathbf{v}_{j}^{\mathsf{T}}\mathbf{\bar{v}}_{j}|$ (7)
where we assumed $\mathbf{v}_{j}^{\mathsf{T}}\mathbf{\bar{v}}_{j}\geq 0$. Note that this assumption does not lose generality because eigenvectors are invariant to change of sign, i.e., if $\mathbf{v}_{j}$ is an eigenvector of $\mathbf{C}$, $-\mathbf{v}_{j}$ is equivalently an eigenvector. Now, we write $\mathbf{v}_{j}-\mathbf{\bar{v}}_{j}=\delta\mathbf{v}_{j}$ as the sum of two components, $\delta\mathbf{v}_{j}=\delta\mathbf{v}_{j\parallel}+\delta\mathbf{v}_{j\perp}$, that are respectively parallel and perpendicular to $\mathbf{v}_{j}$. Due to triangle inequality, we have that $\|\delta\mathbf{v}_{j\perp}\|+\|\delta\mathbf{v}_{j\parallel}\|\leq\|\delta\mathbf{v}_{j\perp}\|\leq\|\delta\mathbf{v}_{j}\|$. Since the columns of $\mathbf{V}$ form an orthonormal basis of $\mathbb{R}^{N}$, the perpendicular component of $\delta\mathbf{v}_{j}$ is the sum of the projections of $\mathbf{\bar{v}}_{j}$ on all columns of $\mathbf{V}$ except $\mathbf{v}_{j}$, i.e., $\delta\mathbf{v}_{j\perp}=\sum_{i=1,i\neq j}^{N}(\mathbf{\bar{v}}_{j}^{\mathsf{T}}\mathbf{v}_{i})\mathbf{v}_{i}$ and the norm becomes $\|\delta\mathbf{v}_{j\perp}\|=\sqrt{\sum_{i=1,i\neq j}^{N}(\mathbf{\bar{v}}_{j}^{\mathsf{T}}\mathbf{v}_{i})^{2}}$. Due to triangle inequality and orthonormality of $\mathbf{V}$, for any $i\neq j$, we have
$\frac{1}{\sqrt{2}}|\mathbf{\bar{v}}_{j}^{\mathsf{T}}\mathbf{v}_{i}|\leq\frac{1}{\sqrt{2}}\|\delta\mathbf{v}_{j\perp}\|\leq\frac{1}{\sqrt{2}}\|\mathbf{v}_{j}-\mathbf{\bar{v}}_{j}\|=\sqrt{1-|\mathbf{v}_{j}^{\mathsf{T}}\mathbf{\bar{v}}_{j}|}.$ (8)
We now provide a general result on VNN stability that will be the starting point to prove Thms. 2 and 3.
@@
$\|\mathbf{H}(\mathbf{\bar{C}})\mathbf{x}-\mathbf{H}(\mathbf{C})\mathbf{x}\|\leq P\sqrt{N}\|\mathbf{E}\|(1+\sqrt{2N})+\mathcal{O}(\|\mathbf{E}\|^{2}).$ (9)
@@
We will make use of the eigendecompositions $\mathbf{C}=\mathbf{V}\mathbf{\Lambda}\mathbf{V}^{\mathsf{T}}$ and $\mathbf{\bar{C}}=\mathbf{\bar{V}}\bar{\mathbf{\Lambda}}\mathbf{\bar{V}}^{\mathsf{T}}$, where $\mathbf{V}$ and $\mathbf{\bar{V}}$ contain in their columns the eigenvectors of $\mathbf{C}$ and $\mathbf{\bar{C}}$, respectively, and $\mathbf{\Lambda}$ and $\bar{\mathbf{\Lambda}}$ contain their eigenvalues on the diagonal. Following (Sihag et al., 2022, eq. (32)-(39)), under the assumption of small perturbation, i.e., $\|\mathbf{E}\|\ll 1$, the stability bound is the sum of three terms:
$\mathbf{H}(\mathbf{\bar{C}})\mathbf{x}-\mathbf{H}(\mathbf{C})\mathbf{x}\approx\underbrace{\sum_{i=0}^{N}\tilde{x}_{i}\sum_{k=0}^{K}h_{k}\sum_{r=0}^{k-1}\mathbf{C}^{r}\lambda_{i}^{k-r-1}(\bar{\lambda}_{i}-\lambda_{i})\mathbf{v}_{i}}_{\textnormal{term 1}}$ (10)
$+\underbrace{\sum_{i=0}^{N}\tilde{x}_{i}\sum_{k=0}^{K}h_{k}\sum_{r=0}^{k-1}\mathbf{C}^{r}\lambda_{i}^{k-r-1}(\lambda_{i}\mathbf{I}_{N}-\mathbf{C})(\mathbf{\bar{v}}_{i}-\mathbf{v}_{i})}_{\textnormal{term 2}}$ (11)
$+\underbrace{\sum_{i=0}^{N}\tilde{x}_{i}\sum_{k=0}^{K}h_{k}\sum_{r=0}^{k-1}\mathbf{C}^{r}\lambda_{i}^{k-r-1}((\bar{\lambda}_{i}-\lambda_{i})\mathbf{I}_{N}-\mathbf{E})(\mathbf{\bar{v}}_{i}-\mathbf{v}_{i})}_{\textnormal{term 3}}$ (12)
@@
where $\tilde{x}_{i}$ is the $i$-th component of the Covariance Fourier Transform of a generic graph signal $\mathbf{x}$, i.e., $\mathbf{\tilde{x}}=\mathbf{V}^{\mathsf{T}}\mathbf{x}$, and we inverted term 1 and 2 compared to (Sihag et al., 2022) for ease of explanation in the following. We now analyze the three terms separately.
Term 1. Leveraging (Sihag et al., 2022, eq. (59)-(61)), we have that
$\sum_{i=0}^{N}\tilde{x}_{i}\sum_{k=0}^{K}h_{k}\sum_{r=0}^{k-1}\mathbf{C}^{r}\lambda_{i}^{k-r-1}(\bar{\lambda}_{i}-\lambda_{i})\mathbf{v}_{i}=\sum_{i=0}^{N}\tilde{x}_{i}h^{\prime}(\lambda_{i})(\bar{\lambda}_{i}-\lambda_{i})\mathbf{v}_{i}$ (13)
@@
where $h^{\prime}(\lambda_{i})$ is the derivative of the frequency response of the filter and is bounded in absolute value by $P$. Using Weyl’s theorem (Golub & van Loan, 2013, Theorem 8.1.6), we have that $|\bar{\lambda}_{i}-\lambda_{i}|\leq\|\mathbf{E}\|$. Therefore, by taking the norm and using the fact that $\sum_{i=1}^{N}|\tilde{x}_{i}|\leq\sqrt{N}\|\mathbf{\tilde{x}}\|=\sqrt{N}\|\mathbf{x}\|\leq\sqrt{N}$, we get
$\left\|\sum_{i=0}^{N}\tilde{x}_{i}h^{\prime}(\lambda_{i})(\bar{\lambda}_{i}-\lambda_{i})\mathbf{v}_{i}\right\|\leq\sum_{i=0}^{N}|\tilde{x}_{i}||h^{\prime}(\lambda_{i})|\|\mathbf{E}\|\|\mathbf{v}_{i}\|\leq$ (14)
$\sum_{i=0}^{N}|\tilde{x}_{i}|P\|\mathbf{E}\|\leq$ (15)
$P\sqrt{N}\|\mathbf{E}\|.$ (16)
Term 2. Following (Sihag et al., 2022, eq. (40)-(54)), the norm of term 2 is bounded by
$\left\|\sum_{i=0}^{N}\tilde{x}_{i}\sum_{k=0}^{K}h_{k}\sum_{r=0}^{k-1}\mathbf{C}^{r}\lambda_{i}^{k-r-1}(\lambda_{i}\mathbf{I}_{N}-\mathbf{C})(\mathbf{\bar{v}}_{i}-\mathbf{v}_{i})\right\|\leq\sqrt{N}\sum_{i=1}^{N}|\tilde{x}_{i}|\max_{j}|h(\lambda_{i})-h(\lambda_{j})||\mathbf{v}_{j}^{\mathsf{T}}\mathbf{\bar{v}}_{i}|$ (17)
@@
where $h(\lambda)$ is the frequency response of the covariance filter. Note that in equation 17 we have a term $\sqrt{N}$ that does not appear in (Sihag et al., 2022, eq. (54)) because we consider operator norm instead of uniform norm. We now leverage the sin-theta theorem (Stewart & guang Sun, 1990, Theorem V.3.6): for $j=1,\dots,N$, we have that
$\frac{1}{2}\|\mathbf{v}_{j}-\mathbf{\bar{v}}_{j}\|^{2}=1-|\mathbf{v}_{j}^{\mathsf{T}}\mathbf{\bar{v}}_{j}|\leq\frac{\|\mathbf{E}\|^{2}}{\min_{i}|\lambda_{i}-\lambda_{i+1}|^{2}}.$ (18)
By using Lemma 2, we obtain:
$|\mathbf{v}_{j}^{\mathsf{T}}\mathbf{\bar{v}}_{i}|\leq\sqrt{2}\sqrt{1-|\mathbf{v}_{j}^{\mathsf{T}}\mathbf{\bar{v}}_{j}|}\leq\frac{\sqrt{2}\|\mathbf{E}\|}{\min_{i}|\lambda_{i}-\lambda_{i+1}|}\quad\forall i\neq j.$ (19)
By plugging equation 19 into equation 17, we have
@@
$\sqrt{2N}\|\mathbf{E}\|\sum_{i=1}^{N}|\tilde{x}_{i}|\max_{j\neq i}\frac{|h(\lambda_{i})-h(\lambda_{j})|}{\min_{i}|\lambda_{i}-\lambda_{i+1}|}\leq$ (21)
$\sqrt{2N}\|\mathbf{E}\|\sum_{i=1}^{N}|\tilde{x}_{i}|\max_{j\neq i}\frac{|h(\lambda_{i})-h(\lambda_{j})|}{|\lambda_{i}-\lambda_{j}|}\leq$ (22)
$\sqrt{2N}\|\mathbf{E}\|\sum_{i=1}^{N}|\tilde{x}_{i}|P\leq$ (23)
$\sqrt{2}N\|\mathbf{E}\|P$ (24)

### v2

@@ VI Related Works
Sparse PCA. The instability of the finite-sample covariance matrix is widely studied for PCA [8, 9, 10] and various regularized estimators have been proposed including covariance shrinkage [30, 50], lasso penalties [31, 51], and thresholding [25, 18]. We leverage here hard and soft thresholding given their benefits in PCA in low-data sparse-covariance settings and study their impact on VNN stability, and we extend the analysis to the generic case of a dense true covariance matrix via stochastic sparsification which is generally not addressed by sparse PCA approaches.
@@
GNNs and VNNs. VNNs are graph convolutional neural networks operating on the covariance graph [11], performing an operation that draws analogies with PCA, but achieves improved stability to finite-sample estimation errors – this has been proved by extending the small perturbation analysis of GNNs [12, 52, 53, 54, 55, 56] to covariance graphs. Due to their robustness, VNNs have shown successful performance in brain data processing [13], transferability to high-dimensional data [14], temporal settings [15] and biased datasets [16]. All these results, however, hold for the dense covariance matrix which is suboptimal in low-data regimes and computationally heavy. Here, we propose S-VNNs and study their stability when the true covariance matrix is sparse and dense. This merits a different treatment than the perturbation assumption studied in [11] and overcomes the limitations of dense VNNs. Our stochastically sparsified VNN draws analogies with the stability of GNNs to random link drops [36, 57] but we generalize those findings to the more challenging case where all sparsified probabilities differ rather than being identical. We also propose design strategies for the sparsified probabilities based on our theoretical analysis.
@@
Graph sparsification in GNNs. Graph sparsification techniques have been used for multiple purposes in GNNs but predominantly to improve: (i) scalability in large and dense graphs [48, 58, 59, 60, 61, 62]; (ii) GNN expressiveness in graph classification tasks and alleviate oversmoothing [63, 64, 65, 66, 67]; and (iii) enhance interpretability [68, 69, 70]. Our S-VNNs aim on the one hand to improve scalability and on the other to preserve the tractability and theoretical links with covariance-based data processing. While our sparsification approaches could be applied also to weighted graphs, we focus on covariance sparsification to characterize theoretically how this affects stability w.r.t. covariance estimation errors and spurious correlations. We also propose a theoretically-grounded stochastic sparsification tailored to sparsity generic dense covariance matrices, which achieves a superior performance than alternatives not developed for this setting.
VII Conclusion
@@
In this work, we develop sparse covariance neural networks (S-VNNs) and study the effects of sparsification techniques on covariance neural networks. We show that S-VNNs can be more stable to finite-data effects on the sample covariance matrix and more computationally efficient. When the true covariance is sparse, we propose S-VNNs with hard/soft thresholding and show that they transfer to the nominal VNN with a rate inversely proportional to the square root of the number of data points. When the true covariance is dense, we put forth a stochastic sparsified approach to reduce the computational complexity while maintaining mathematical tractability and stability. In the framework of stochastic sparsification, we further propose principled strategies, i.e., ACV and RCV, that allow to tune sparsification impact and computation efficiency. Experimental results on both synthetic and real datasets show that the proposed S-VNNs improve substantially the computation time and achieve competitive or better performance compared with nominal VNNs, thus validating that spurious correlations are present and damaging in datasets. Future work will investigate other potential covariance estimators that can be used in S-VNNs and characterize the regularization effect of stochastic sparsification, which allows the model to observe multiple instances of the sparsified covariance with potential benefits for robustness, expressiveness, generalization and reduced oversmoothing similarly to dropout-like techniques and graph sparsification methods [63, 71, 64].
Appendix A Proof of Lemma˜1
@@
$\begin{split}\left\|\mathbf{V}^{\mathsf{T}}\mathbf{x}-\mathbf{\hat{V}}^{\mathsf{T}}\mathbf{x}\right\|&\leq\|\mathbf{V}-\mathbf{\hat{V}}\|\|\mathbf{x}\|\\ &\leq\sqrt{N}\max_{i}\left\|\mathbf{v}_{i}-\mathbf{\hat{v}}_{i}\right\|\end{split}$ (5)
where we used spectral norm properties. Using the sin-theta theorem [72, Theorem V.3.6], we have that
$\max_{i}\left\|\mathbf{v}_{i}-\mathbf{\hat{v}}_{i}\right\|\leq\frac{\sqrt{2}\|\mathbf{E}\|}{\min_{j}|\lambda_{j}-\lambda_{j+1}|}$ (6)
where $\|\mathbf{E}\|=\|\mathbf{C}-\mathbf{\hat{C}}\|$. We now leverage the result from [24, Theorem 5.6.1] which shows that $\|\mathbf{E}\|\leq\mathcal{O}(t^{-1/2})$ with probability 1o(1)1-o(1), and this concludes the proof.
Appendix B Proof of Thms.˜3 and 4
B-A Preliminaries
@@
Consider the true and the thresholded sample covariance matrices with their respective eigendecompositions $\mathbf{C}=\mathbf{V}\mathbf{\Lambda}\mathbf{V}^{\mathsf{T}}$ and $\mathbf{\bar{C}}=\mathbf{\bar{V}}\bar{\mathbf{\Lambda}}\mathbf{\bar{V}}^{\mathsf{T}}$ and let $\mathbf{E}=\mathbf{C}-\mathbf{\bar{C}}$ be the error estimation matrix such that $\|\mathbf{E}\|\ll 1$ (i.e., small perturbation). The following holds:
$|\mathbf{v}_{j}^{\mathsf{T}}\mathbf{\bar{v}}_{i}|\leq\frac{\|\mathbf{E}\|}{|\lambda_{i}-\lambda_{j}|}+\mathcal{O}(\|\mathbf{E}\|^{2})$ (7)
@@
From [73, Chapter 2, (10.2)], using the assumptions in Thms.˜3 and 4 of $\mathbf{C}$ being positive definite with distinct eigenvalues, we have that the first-order approximation of the $i$-th estimated covariance eigenvector $\mathbf{\tilde{v}}_{i}$ is
$\mathbf{\bar{v}}_{i}\approx\mathbf{v}_{i}+\sum_{k=1,k\neq i}^{N}\frac{\mathbf{v}_{k}^{\mathsf{T}}\mathbf{E}\mathbf{v}_{i}}{\lambda_{k}-\lambda_{i}}+\mathbf{D}$ (8)
with $\|\mathbf{D}\|=\mathcal{O}(\|\mathbf{E}\|^{2})$. By projecting on $\mathbf{v}_{j}$ and taking the absolute value, we get
$|\mathbf{v}_{j}^{\mathsf{T}}\mathbf{\bar{v}}_{i}|\leq\frac{|\mathbf{v}_{j}^{\mathsf{T}}\mathbf{E}\mathbf{v}_{i}|}{|\lambda_{i}-\lambda_{j}|}+\mathcal{O}(\|\mathbf{E}\|^{2})\leq$ (9)
$\frac{\|\mathbf{E}\|}{|\lambda_{i}-\lambda_{j}|}+\mathcal{O}(\|\mathbf{E}\|^{2})$ (10)
where the last step follows from the definition of spectral norm: $\|\mathbf{E}\|=\max_{\mathbf{v},\|\mathbf{v}\|_{2}=1}|\mathbf{v}^{\mathsf{T}}\mathbf{E}\mathbf{v}|$. ∎
We now provide a general result on VNN stability that will be the starting point to prove Thms.˜3 and 4.
@@
$\|\mathbf{H}(\mathbf{\bar{C}})\mathbf{x}-\mathbf{H}(\mathbf{C})\mathbf{x}\|\leq P\sqrt{N}\|\mathbf{E}\|(1+\sqrt{N})+\mathcal{O}(\|\mathbf{E}\|^{2}).$ (11)
@@
We will make use of the eigendecompositions $\mathbf{C}=\mathbf{V}\mathbf{\Lambda}\mathbf{V}^{\mathsf{T}}$ and $\mathbf{\bar{C}}=\mathbf{\bar{V}}\bar{\mathbf{\Lambda}}\mathbf{\bar{V}}^{\mathsf{T}}$, where $\mathbf{V}$ and $\mathbf{\bar{V}}$ contain in their columns the eigenvectors of $\mathbf{C}$ and $\mathbf{\bar{C}}$, respectively, and $\mathbf{\Lambda}$ and $\bar{\mathbf{\Lambda}}$ contain their eigenvalues on the diagonal. Following [11, eq. (32)-(39)], under the assumption of small perturbation, i.e., $\|\mathbf{E}\|\ll 1$, the stability bound is the sum of three terms:
$\mathbf{H}(\mathbf{\bar{C}})\mathbf{x}-\mathbf{H}(\mathbf{C})\mathbf{x}\approx\underbrace{\sum_{i=1}^{N}\tilde{x}_{i}\sum_{k=0}^{K}h_{k}\sum_{r=0}^{k-1}\mathbf{C}^{r}\lambda_{i}^{k-r-1}(\bar{\lambda}_{i}-\lambda_{i})\mathbf{v}_{i}}_{\textnormal{term 1}}$ (12)
$+\underbrace{\sum_{i=1}^{N}\tilde{x}_{i}\sum_{k=0}^{K}h_{k}\sum_{r=0}^{k-1}\mathbf{C}^{r}\lambda_{i}^{k-r-1}(\lambda_{i}\mathbf{I}_{N}-\mathbf{C})(\mathbf{\bar{v}}_{i}-\mathbf{v}_{i})}_{\textnormal{term 2}}$ (13)
$+\underbrace{\sum_{i=1}^{N}\tilde{x}_{i}\sum_{k=0}^{K}h_{k}\sum_{r=0}^{k-1}\mathbf{C}^{r}\lambda_{i}^{k-r-1}((\bar{\lambda}_{i}-\lambda_{i})\mathbf{I}_{N}-\mathbf{E})(\mathbf{\bar{v}}_{i}-\mathbf{v}_{i})}_{\textnormal{term 3}}$ (14)
@@
where $\tilde{x}_{i}$ is the $i$-th component of the Covariance Fourier Transform of a generic graph signal $\mathbf{x}$, i.e., $\mathbf{\tilde{x}}=\mathbf{V}^{\mathsf{T}}\mathbf{x}$, and we inverted term 1 and 2 compared to [11] for ease of explanation in the following. We now analyze the three terms separately.
Term 1. Leveraging [11, eq. (59)-(61)], we have that
$\sum_{i=1}^{N}\tilde{x}_{i}\sum_{k=0}^{K}h_{k}\sum_{r=0}^{k-1}\mathbf{C}^{r}\lambda_{i}^{k-r-1}(\bar{\lambda}_{i}-\lambda_{i})\mathbf{v}_{i}=$ (15)
$\sum_{i=1}^{N}\tilde{x}_{i}h^{\prime}(\lambda_{i})(\bar{\lambda}_{i}-\lambda_{i})\mathbf{v}_{i}$ (16)
@@
where $h^{\prime}(\lambda_{i})$ is the derivative of the frequency response of the filter and is bounded in absolute value by $P$. Using Weyl’s theorem [74, Theorem 8.1.6], we have that $|\bar{\lambda}_{i}-\lambda_{i}|\leq\|\mathbf{E}\|$. Therefore, by taking the norm and using the fact that $\sum_{i=1}^{N}|\tilde{x}_{i}|\leq\sqrt{N}\|\mathbf{\tilde{x}}\|=\sqrt{N}\|\mathbf{x}\|\leq\sqrt{N}$, we get
$\left\|\sum_{i=1}^{N}\tilde{x}_{i}h^{\prime}(\lambda_{i})(\bar{\lambda}_{i}-\lambda_{i})\mathbf{v}_{i}\right\|\leq\sum_{i=1}^{N}|\tilde{x}_{i}||h^{\prime}(\lambda_{i})|\|\mathbf{E}\|\|\mathbf{v}_{i}\|\leq$ (17)
$\sum_{i=1}^{N}|\tilde{x}_{i}|P\|\mathbf{E}\|\leq P\sqrt{N}\|\mathbf{E}\|.$ (18)
Term 2. Following [11, eq. (40)-(54)], the norm of term 2 is bounded by
$\left\|\sum_{i=1}^{N}\tilde{x}_{i}\sum_{k=0}^{K}h_{k}\sum_{r=0}^{k-1}\mathbf{C}^{r}\lambda_{i}^{k-r-1}(\lambda_{i}\mathbf{I}_{N}-\mathbf{C})(\mathbf{\bar{v}}_{i}-\mathbf{v}_{i})\right\|\leq$ (19)
$\sqrt{N}\sum_{i=1}^{N}|\tilde{x}_{i}|\max_{j}|h(\lambda_{i})-h(\lambda_{j})||\mathbf{v}_{j}^{\mathsf{T}}\mathbf{\bar{v}}_{i}|$ (20)
@@
where $h(\lambda)$ is the frequency response of the covariance filter. Note that in (19) we have a term $\sqrt{N}$ that does not appear in [11, eq. (54)] because we consider operator norm instead of uniform norm.
Using Lemma˜2 in (20), we get
@@
$\sqrt{N}\|\mathbf{E}\|\sum_{i=1}^{N}|\tilde{x}_{i}|\max_{j\neq i}\frac{|h(\lambda_{i})-h(\lambda_{j})|}{|\lambda_{i}-\lambda_{j}|}\leq$ (22)
$\sqrt{N}\|\mathbf{E}\|\sum_{i=1}^{N}|\tilde{x}_{i}|P\leq N\|\mathbf{E}\|P$ (23)